i'm the middle number of three consecutive numbers.the product of the three consecutive numbers is 17550
Answer by stanbon(75887)

i'm the middle number of three consecutive numbers.the product of the three consecutive numbers is 17550
----------------
1st: x-1
2nd: x
3rd: x+1
------
Equation:
x(x^2-1) = 17550
x^3-x-17550 = 0
-----
x = 26
x-1 = 25
x+1 = 27
---------------
Cheers,
Stan H.
================